# Conditional entropy and information of quantum processes

### v1

@@ V.2 Geometric Rényi conditional entropies
$\widehat{{S}}_{\alpha}[A|B]_{\mathcal{N}}$ $=\log|A^{\prime}|-\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty},$ (79)
@@
The proof follows by invoking  [34, Lemma 5] in Definition Eq.  (78):
@@
$=\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},|A^{\prime}|\mathbbm{1}_{R_{A}}\otimes\mathbbm{1}_{A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty}$
$=-\log|A^{\prime}|+\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty}.$ (80)
Remark 3.
The geometric Rényi conditional entropy $\widehat{S}_{\alpha}$, for $\alpha\in(1,2]$, of an arbitrary finite-dimensional bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ has a semidefinite representation and can be efficiently computed using a semidefinite programme. It follows by applying [34, Lemma 9] to the r.h.s. of the given identity,
$\log|A^{\prime}|-\widehat{{S}}_{\alpha}[A|B]_{\mathcal{N}}=\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty}.$ (81)
@@ V.3 Conditional min and max entropies
The max-relative entropy between a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, when $|A^{\prime}|<\infty$, and a completely positive map $\mathcal{M}_{A^{\prime}\to A}$ is given as [31]
@@
${S}_{\min}[A|B]_{\mathcal{N}}=\log|A^{\prime}|-\log\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\norm{\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}}_{\infty},$ (85)
@@
$=$ $\log|A^{\prime}|-\log\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\norm{\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}}_{\infty}.$ (86)
@@
where $D_{\min}[\mathcal{N}\parallel\mathcal{M}]:=\sup_{\psi_{RA^{\prime}}\in\mathrm{St}(RA^{\prime})}D_{\min}(\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})\parallel\operatorname{id}_{R}\otimes\mathcal{M}_{A^{\prime}\to A}(\psi_{RA^{\prime}}))$ is the min-relative entropy between completely positive maps $\mathcal{N},\mathcal{M}$.
VI Bipartite processes: Signalling vs Semicausal
@@
if and only if $\mathcal{N}$ is no-signalling from $A^{\prime}$ to $B$, i.e., $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\in\mathrm{NS}_{A^{\prime}\not\to B}$.
If a bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is signalling from $A^{\prime}$ to $B$, i.e., there is causal influence from $A^{\prime}$ to $B$ ($\mathcal{N}\notin{\mathrm{N}S}_{A^{\prime}\not\to B})$, then
@@
Here, $\rho_{RAB}=(\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}B^{\prime}\to AB})(\psi_{RA^{\prime}B^{\prime}})$. In the third equality, we have used $D(\alpha_{XY}\parallel\mathbbm{1}_{X}\otimes\beta_{Y})=D(\alpha_{XY}\parallel\mathbbm{1}_{X}\otimes\alpha_{Y})+D(\alpha_{Y}\parallel\beta_{Y})$ for all $\alpha\in\mathrm{St}(XY)$, $\beta\in\mathrm{St}(Y)$. Further we have invoked Eq. (88) to write $\rho_{RB}=(\operatorname{id}_{R}\otimes\widetilde{\mathcal{N}})(\psi_{RB^{\prime}})$. To obtain the fourth equality we note:
@@
$=\sup_{\psi_{RA^{\prime}B^{\prime}}\in\mathrm{St}(RA^{\prime}B^{\prime})}D\left(\rho_{RAB}\parallel\mathbbm{1}_{A}\otimes\rho_{RB}\right)$ (95)
@@
For Eq. (91), we start from the fourth equality in the above chain of equations and note that $D(\rho_{RAB}\parallel\mathbbm{1}_{A}\otimes\rho_{RB})=-S(A|RB)_{\rho}$. Since the chain of equations is guaranteed to be saturated, i.e., equalities hold, we have an “if and only if" statement. Applying Proposition 2, we arrive at inequality (92). ∎
@@
Remark 4 (Generalized conditional NS-entropy).
Let $S^{\not\to}[A|B]_{\mathcal{N}}$ denote the function in the r.h.s. of Eq. (90) for an arbitrary bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$. This is the candidate A in Eq. (38). Its generalized form can be denoted as $\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$,
$\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}:=\inf_{\mathcal{M}\in{\mathrm{Q}}(B^{\prime},B)}\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}].$ (98)
$\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$ satisfies all the properties in Theorem 2. Therefore, it can be deemed a valid conditional entropy of a bipartite channel. We refer to $\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$ is the von Neumann condition NS-entropy. However, it is not our first preference for the definition of the conditional entropy when compared to $\mathbf{S}[A|B]_{\mathcal{N}}$. $S^{\not\to}[A|B]_{\mathcal{N}}=\inf_{\psi\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}S(A|RB)_{\mathcal{N}(\psi)}$, suffices to optimize over pure states $\psi_{RA^{\prime}B^{\prime}}$, for any channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ and the dimensional bounds
$-\log|A|\leq S^{\not\to}[A|B]_{\mathcal{N}}\leq\log|A|$ (99)
resembles the von Neumann conditional entropy of quantum states and the von Neumann entropy of channel, see Table 1. It does not appear to have the potential for the kind of novelty we desire.
@@
[56] A bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is $k$-extendible with respect to the system $A^{\prime},A$, if there exists a quantum channel $\mathcal{M}\in Q(A^{\prime}_{1}A^{\prime}_{2}\cdots A^{\prime}_{k}B^{\prime},A_{1}A_{2}\cdots A_{k}B)$, with $A^{\prime}{\simeq}A^{\prime}_{1}{\simeq}A^{\prime}_{2}{\simeq\cdots\simeq}A^{\prime}_{k}$ and $A{\simeq}A_{1}{\simeq}A_{2}{\simeq\cdots\simeq}A_{k}$, if the following holds:

### v2

@@ V.2 Geometric Rényi conditional entropies
$\widehat{{S}}_{\alpha}[A|B]_{\mathcal{N}}$ $=-\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty},$ (79)
@@
The proof follows by invoking  [37, Lemma 5] in Definition Eq.  (78):
@@
$=\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\frac{1}{\alpha-1}\log\norm{\operatorname{tr}_{AB}G_{1-\alpha}\left(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B},\mathbbm{1}_{R_{A}}\otimes\mathbbm{1}_{A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}\right)}_{\infty}.$ (80)
Due to [37, Lemma 9], the geometric Rényi conditional entropy $\widehat{S}_{\alpha}$, for $\alpha\in(1,2]$, as in Eq. (80), of an arbitrary finite-dimensional bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ has a semidefinite representation as shown in the following corollary.
Corollary 3.
For $\alpha=1+2^{-\ell}$ with $\ell\in\mathbb{N}$, the geometric Rényi conditional entropy $\widehat{{S}}_{\alpha}[A|B]_{\mathcal{N}}$ takes the form $\widehat{{S}}_{\alpha}[A|B]_{\mathcal{N}}=-2^{\ell}\log y^{*}$, where $y^{*}$ is obtained from the following semidefinite programme:
$\mathrm{minimize:}\ y$
$\mathrm{subject\ to:}\ P\in\mathsf{Herm}(R_{A}AR_{B}B),\{Q_{i}\in\mathsf{Herm}(R_{A}AR_{B}B)\}_{i=0}^{\ell},$
$\qquad\qquad\quad\begin{pmatrix}P&\Gamma^{\mathcal{N}}\\ \Gamma^{\mathcal{N}}&Q_{\ell}\\ \end{pmatrix}\geq 0,\ \bigg{\{}\begin{pmatrix}\Gamma^{\mathcal{N}}&Q_{i}\\ Q_{i}&Q_{i-1}\\ \end{pmatrix}\geq 0\bigg{\}}_{i=1}^{\ell},\ \Gamma^{\mathcal{M}}_{R_{B}B}\geq 0,\ \operatorname{tr}_{B}\Gamma^{\mathcal{M}}=\mathbbm{1}_{R_{B}},$
$\qquad\qquad\quad Q_{0}=\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B},\ \operatorname{tr}_{AB}P\leq y\mathbbm{1}_{R_{A}R_{B}},$ (81)
where $\mathsf{Herm}(R_{A}AR_{B}B)$ is the set of all Hermitian operators in $\mathrm{L}(R_{A}AR_{B}B)$.
@@ V.3 Conditional min and max entropies
The max-relative entropy between a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, when $|A^{\prime}|<\infty$, and a completely positive map $\mathcal{M}\in{\mathrm{C}P}(A^{\prime},A)$ is given as [55, Lemma 12]
@@
${S}_{\min}[A|B]_{\mathcal{N}}=-\log\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\norm{\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}}_{\infty},$ (85)
@@
$=$ $-\log\min_{\mathcal{M}\in\mathrm{Q}(B^{\prime},B)}\norm{\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\mathbbm{1}_{R_{A}A}\otimes(\Gamma^{\mathcal{M}}_{R_{B}B})^{-\frac{1}{2}}}_{\infty}.$ (86)
The conditional min-entropy of a bipartite channel has a semidefinite representation, as shown in the following corollary.
Corollary 4.
The conditional min-entropy can be estimated via the following semidefinite program:
$S_{\min}[A|B]_{\mathcal{N}}=-\log\min\bigg{\{}\frac{\operatorname{tr}(\widetilde{M})}{\absolutevalue{B^{\prime}}}\ |\ \Gamma^{\mathcal{N}}\leq\mathbbm{1}_{R_{A}A}\otimes\widetilde{M}_{R_{B}B},\ \widetilde{M}\geq 0,\ \operatorname{tr}_{B}\widetilde{M}=\operatorname{tr}(\widetilde{M})\frac{\mathbbm{1}_{R_{B}}}{\absolutevalue{B^{\prime}}}\bigg{\}}.$ (87)
Proof.
As a consequence of Eq. (82) and Eq. (84), we have
$S_{\min}[A|B]_{\mathcal{N}}$ $=-\inf_{\mathcal{M}\in Q(B^{\prime},B)}D_{\max}(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\parallel\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}).$ (88)
From the semidefinite representation of the max-relative entropy as shown in Eq. (18), we have
$\inf_{\mathcal{M}\in Q(B^{\prime},B)}D_{\max}(\Gamma^{\mathcal{N}}_{R_{A}AR_{B}B}\parallel\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B})$
$=\log\min\{y\ |\ \Gamma^{\mathcal{N}}\leq y(\mathbbm{1}_{R_{A}A}\otimes\Gamma^{\mathcal{M}}_{R_{B}B}),\ \Gamma^{\mathcal{M}}\geq 0,\ \operatorname{tr}_{B}\Gamma^{\mathcal{M}}=\mathbbm{1}_{R_{B}}\}.$ (89)
With this we prove the corollary by defining $\widetilde{M}:=y\Gamma^{\mathcal{M}}$ and noting $y=\operatorname{tr}(\widetilde{M})/\absolutevalue{B^{\prime}}$ |. ∎
Figure 2: We plot the conditional entropies $\widehat{S}_{\alpha}[A|B]_{\mathcal{S}^{p,U}}$ and ${S}_{\min}[A|B]_{\mathcal{S}^{p,U}}$ of bipartite channels $\mathcal{S}^{p,U}$ for $U\in\{{\mathrm{S}WAP,CNOT}\}$ vs channels’ mixture parameter $p\in[0,1]$. The dashed and dashed-dotted and dotted lines represent the geometric Rényi conditional entropies with $\alpha=1.0625$, $\alpha=1.25$, and $\alpha=2$, respectively, all of which are estimated using Corollary 3. The solid lines represent the conditional min-entropy $S_{\min}[A|B]_{\mathcal{S}^{p,U}}$ estimated using Corollary 4. Fig.(a) represents the mixture of SWAP and the identity gates (Eq. (90)) and Fig.(b) represents the mixture of CNOT and the identity gates (Eq. (91)).
Using the semidefinite representations in Corollaries 3 and  4, we estimate the geometric Rényi conditional entropy and the conditional min-entropy of the probabilistic mixtures of the two-qubit SWAP and identity gates and the two-qubit CNOT and identity gates:
$\mathcal{S}^{p,\mathrm{SWAP}}_{A^{\prime}B^{\prime}\to AB}(\cdot)=p\mathrm{SWAP}(\cdot)\mathrm{SWAP}^{\dagger}+(1-p)\operatorname{id}(\cdot),\quad p\in[0,1],$ (90)
$\mathcal{S}^{p,\mathrm{CNOT}}_{A^{\prime}B^{\prime}\to AB}(\cdot)=p\mathrm{CNOT}(\cdot)\mathrm{CNOT}^{\dagger}+(1-p)\operatorname{id}(\cdot),\quad p\in[0,1].$ (91)
In Fig. 2, we plot the conditional entropies $\widehat{S}_{\alpha}[A|B]_{\mathcal{S}^{p,U}}$ and ${S}_{\min}[A|B]_{\mathcal{S}^{p,U}}$ of the channels in Eqs. (90) and (91) with respect to the channels’ mixing parameter $p$. For the case of the geometric Rényi conditional entropies $\widehat{S}_{\alpha}[A|B]_{\mathcal{S}^{p,U}}$ with $U\in\{{\mathrm{S}WAP,CNOT}\}$, we choose $\alpha=2$, $\alpha=1.25$, and $\alpha=1.0625$, which correspond to $\ell=0$, $\ell=2$ and $\ell=4$, respectively, as in Corollary 3. Fig. 2(a) plots for the mixture of SWAP and the identity (Eq. (90)) and Fig. 2(b) plots for the mixture of CNOT and the identity (Eq. (91)). The plots show $S_{\min}[A|B]_{\mathcal{S}^{p,U}}\leq\widehat{S}_{\alpha_{1}}[A|B]_{\mathcal{S}^{p,U}}\leq\widehat{S}_{\alpha_{2}}[A|B]_{\mathcal{S}^{p,U}}$ for $\alpha_{1}\geq\alpha_{2}$ in accordance with Eq. (45).
@@
where $D_{\min}[\mathcal{N}\parallel\mathcal{M}]:=\sup_{\psi_{RA^{\prime}}\in\mathrm{St}(RA^{\prime})}D_{\min}(\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})\parallel\operatorname{id}_{R}\otimes\mathcal{M}_{A^{\prime}\to A}(\psi_{RA^{\prime}}))$ is the min-relative entropy between $\mathcal{N},\mathcal{M}\in{\mathrm{C}P}(A^{\prime},A)$.
VI Bipartite processes: Signaling vs Semicausal
@@
if and only if $\mathcal{N}$ is no-signaling from $A^{\prime}$ to $B$, i.e., $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\in\mathrm{NS}_{A^{\prime}\not\to B}$.
A bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is signaling from $A^{\prime}$ to $B$, i.e., there is causal influence from $A^{\prime}$ to $B$ ($\mathcal{N}\notin{\mathrm{N}S}_{A^{\prime}\not\to B})$, if and only if
@@
Here, $\rho_{RAB}:=(\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}B^{\prime}\to AB})(\psi_{RA^{\prime}B^{\prime}})$. In the third equality, we have used $D(\alpha_{XY}\parallel\mathbbm{1}_{X}\otimes\beta_{Y})=D(\alpha_{XY}\parallel\mathbbm{1}_{X}\otimes\alpha_{Y})+D(\alpha_{Y}\parallel\beta_{Y})$ for all $\alpha\in\mathrm{St}(XY)$, $\beta\in\mathrm{St}(Y)$. Further we have invoked Eq. (93) to write $\rho_{RB}=(\operatorname{id}_{R}\otimes\widetilde{\mathcal{N}})(\psi_{RB^{\prime}})$. To obtain the fourth equality we note:
@@
For Eq. (96), we start from the fourth equality in the above chain of equations and note that $D(\rho_{RAB}\parallel\mathbbm{1}_{A}\otimes\rho_{RB})=-S(A|RB)_{\rho}$. Since the chain of equations is guaranteed to be saturated, i.e., equalities hold, we have an “if and only if" statement. We now apply Proposition 2 to arrive at inequality (97). ∎
@@
[59] A bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is $k$-extendible for a natural number $k\geq 2$ with respect to the system $A^{\prime},A$, if there exists a quantum channel $\mathcal{M}\in{\mathrm{Q}}(A^{\prime}_{1}A^{\prime}_{2}\cdots A^{\prime}_{k}B^{\prime},A_{1}A_{2}\cdots A_{k}B)$, where $A^{\prime}{\simeq}A^{\prime}_{1}{\simeq}A^{\prime}_{2}{\simeq\cdots\simeq}A^{\prime}_{k}$ and $A{\simeq}A_{1}{\simeq}A_{2}{\simeq\cdots\simeq}A_{k}$, having the following properties:
@@
VI.1 Generalized conditional NS-entropy
Generalized conditional NS-entropy.— Let $S^{\not\to}[A|B]_{\mathcal{N}}$ denote the function in the r.h.s. of Eq. (95) for an arbitrary bipartite quantum channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$. This is the candidate A in Eq. (40). Its generalized form can be denoted as $\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$,
$\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}:=\mathbf{D}[\mathcal{N}_{A^{\prime}B^{\prime}\to AB}\|\mathcal{R}_{A\to A}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}].$ (107)
$\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$ satisfies all the properties in Theorem 2. Therefore, it can be deemed a valid conditional entropy of a bipartite channel. We refer to $\mathbf{S}^{\not\to}[A|B]_{\mathcal{N}}$ as the generalized conditional NS-entropy, where NS denotes connection to no-signaling. However, it is not our first preference for the definition of the conditional entropy when compared to $\mathbf{S}[A|B]_{\mathcal{N}}$.
For the von Neumann conditional NS-entropy $S^{\not\to}[A|B]_{\mathcal{N}}$ of a bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, we have $S^{\not\to}[A|B]_{\mathcal{N}}=\inf_{\psi\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}S(A|RB)_{\mathcal{N}(\psi)}$, where it suffices to optimize over pure states $\psi_{RA^{\prime}B^{\prime}}$; the dimensional bounds
$-\log|A|\leq S^{\not\to}[A|B]_{\mathcal{N}}\leq\log|A|$ (108)
resemble that of the von Neumann conditional entropy of quantum states and the von Neumann entropy of a channel $\mathcal{N}_{A^{\prime}B^{\prime}\to A}=\operatorname{tr}_{B}\circ\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, see Table 1.
Let $\mathcal{V}^{\mathcal{N}}_{A^{\prime}B^{\prime}\to ABE}$ denote an isometric extension channel of $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$, i.e.,
$\operatorname{tr}_{E}\circ\mathcal{V}^{\mathcal{N}}_{A^{\prime}B^{\prime}\to ABE}=\mathcal{N}_{A^{\prime}B^{\prime}\to AB}.$ (109)
A complementary channel $\mathcal{N}^{c}_{A^{\prime}B^{\prime}\to E}$ of a bipartite channel $\mathcal{N}_{A^{\prime}B^{\prime}\to AB}$ is given as
$\mathcal{N}^{c}_{A^{\prime}B^{\prime}\to E}=\operatorname{tr}_{AB}\circ\mathcal{V}^{\mathcal{N}}_{A^{\prime}B^{\prime}\to ABE}.$ (110)
The von Neumann conditional NS-entropy obeys the following relation,
${S}^{\not\to}[A|B]_{\mathcal{N}}+\sup_{\rho\in{\mathrm{S}t}(A^{\prime}B^{\prime})}S(A|E)_{\mathcal{V}^{\mathcal{N}}(\rho)}=0.$ (111)
It follows because of the duality relation of the von Neumann conditional entropy, $S(A|B)_{\phi}+S(A|E)_{\phi}=0$ for any pure state $\phi_{ABE}$. Then,
$\inf_{\outerproduct{\psi}{\psi}\in{\mathrm{S}t}(RA^{\prime}B^{\prime})}S(A|RB)_{\mathcal{N}(\outerproduct{\psi}{\psi})}=\inf_{\rho\in{\mathrm{S}t}(A^{\prime}B^{\prime})}-S(A|E)_{\mathcal{V}^{\mathcal{N}}(\rho)}=-\sup_{\rho\in{\mathrm{S}t}(A^{\prime}B^{\prime})}S(A|E)_{\mathcal{V}^{\mathcal{N}}(\rho)}.$ (112)